OpenStudy (anonymous):

A 1200kg safe is 1.6m above a heavy-duty spring when the rope holding the safe breaks. The safe hits the spring and compresses it 50cm . Find k

OpenStudy (gorv):

F=kx

OpenStudy (gorv):

k = spring constant x= change in length

OpenStudy (gorv):

f=m*g

OpenStudy (gorv):

can u do it now??

OpenStudy (gorv):

m=12000 f=9.8 x=50/100=0.5 m

OpenStudy (gorv):

divide both side by x

OpenStudy (gorv):

f/x=k

OpenStudy (anonymous):

23520 N/m is that right?